# PreprintToPaper dataset: connecting bioRxiv preprints with journal publications

### v1

@@ Abstract
The PreprintToPaper dataset connects bioRxiv preprints with their corresponding journal publications, enabling large-scale analysis of the preprint-to-publication process. It comprises metadata for 145,517 preprints from two periods, 2016–2018 (pre-pandemic) and 2020–2022 (pandemic), retrieved via the bioRxiv and Crossref APIs. Each record includes bibliographic information such as titles, abstracts, authors, institutions, submission dates, licenses, and subject categories, alongside enriched publication metadata including journal names, publication dates, author lists, and further information. Preprints are categorized into three groups: Published (formally linked to a journal article), Preprint Only (unpublished), and Gray Zone (potentially published but unlinked). To enhance reliability, title and author similarity scores were calculated, and a human-annotated subset of 299 records was created for evaluation of Gray Zone cases. The dataset supports diverse applications, including studies of scholarly communication, open science policies, bibliometric tool development, and natural language processing research on textual changes between preprints and their published versions. The dataset is publicly available in CSV format via Zenodo.
@@
The PreprintToPaper dataset [3], which we describe in this paper, links bioRxiv preprints with their subsequent journal publications, allowing for large-scale analysis of the preprint-to-publication process. It includes metadata on more than 145,000 preprints from two distinct periods (2016–2018, the pre-pandemic period, and 2020–2022, the COVID-19 pandemic period), with information on titles, authors, abstracts, institutions, licenses, submission and publication dates, and subject categories. Preprints are categorized into “Published”, “Preprint Only”, and “Gray Zone” cases, based on automated linking procedures combining DOI information, title similarity, and author similarity. In addition to the main dataset, a human-annotated subset of gray zone cases is provided to support evaluation of the automated matching approach  [^1] This subset includes records with title similarity (title_match_score = 0.75) and was independently checked by two annotators. Further details on the annotation procedure is provided in the Methods section (Step 5: Gray Zone Verification)..
The motivation for creating the PreprintToPaper dataset is to provide a comprehensive resource for studying how preprints evolve into published journal articles—or remain unpublished [4, 5]. While some preprints are eventually linked to formal publications, many others stay as preprints without ever appearing in journals. This dataset is the first systematic effort to automatically collect and link metadata from bioRxiv [6] with publication records, making it possible to analyze patterns in the preprint-to-publication process at scale [7].
@@
  • Natural language processing and text analysis: Supports research on linguistic and structural changes between preprints and their published versions, such as title reformulations, abstract rewriting, and shifts in author order or contribution statements.
@@
Research on bioRxiv has shown that articles posted as preprints tend to receive more citations and online attention than comparable articles not deposited, although the causal mechanisms behind this effect remain uncertain [12]. Surveys of authors indicate that the main motivations for posting preprints are to increase visibility and accelerate dissemination, while concerns about peer review and lack of awareness are the main barriers [2]. Large-scale bibliometric analyses further confirm a citation and altmetric advantage for bioRxiv-linked publications, with preprints themselves increasingly cited and shared on social media [12]. During the COVID-19 pandemic, preprints on bioRxiv and medRxiv played an unprecedented role, being accessed, cited, and shared at higher rates, and influencing both journalistic reporting and policy discussions [1]. An analysis of biomedical publishing during the first months of the COVID-19 pandemic reveals a sharp rise in publication volume, faster acceptance times for COVID-19 papers, and reduced international collaboration, with these shifts occurring largely at the expense of non-COVID-19 research [8]. Other studies of bioRxiv preprints find that around 30% remain unpublished, most are posted too late to enable feedback, and nearly half of published preprints are concentrated in four major publishers, suggesting the platform is used more for priority setting and visibility than for prepublication peer review [4, 13].
@@
The PreprintToPaper dataset offers two main advantages over the covid19_preprints dataset and the Rxivist database:
  • It stores both initial and final versions of preprints, as well as detailed metadata (title, abstract, authors, journal, publication date) about their published versions.
  • It identifies articles that are not formally linked but may have been published (Gray Zone cases).
Find a detailed comparison of the three datasets in Table 1.
Table 1: Comparison of the covid19_preprints, Rxivist, and PreprintToPaper datasets
Aspect covid19_preprints Rxivist PreprintToPaper (Ours)
Time span 2017–2021 Nov 2013 – Feb 2023 2016–2018 (pre-pandemic) and 2020–2022 (pandemic)
Coverage Multiple preprint servers (bioRxiv, arXiv, OSF, ESSOAR, Figshare, Zenodo, etc.) bioRxiv and medRxiv bioRxiv
Scope COVID-19 related preprints only All subjects (bioRxiv + medRxiv) All subjects (bioRxiv)
Size 66,130 records 223,541 records 145,517 records
Preprint information Source, identifier (DOI), posted date, title, abstract Title, abstract, DOI, category, submission date, authors, institutions Title, abstract (initial + last), DOI, category, submission date (initial + last), authors (initial + last), corresponding authors, institutions, license, version number, category, fulltext JATS XML
Publication information None Published DOI, journal name Published DOI, title, abstract, authors, journal name, publication date (from Crossref)
Versioning Only one version per preprint Only one version per preprint Both the first and last versions at the relevant period
Publication linking None Published DOIs partially incomplete (crawling stopped early) Official bioRxiv links + enriched with Crossref + additional Gray Zone investigation ($\approx$19,000 potential unlinked publications)
@@ Methods
Figure 1 illustrates the data generation workflow for the PreprintToPaper dataset. The process consists of six main steps: (1) collecting metadata from the bioRxiv preprint platform; (2) retrieving additional metadata from Crossref based on the published DOIs listed in the bioRxiv records; (3) categorizing preprints into published and preprint-only; (4) identifying unlinked publications in bioRxiv (Gray Zone); (5) verifying Gray Zone preprints using author-based metrics (author match score and author count difference); and (6) generating the final dataset.
Figure 1: PreprintToPaper dataset creation workflow. In Step 5 (Gray Zone verification), a human-annotated subset of 299 rows was created, limited to cases with a title match score of 0.75 and covering both periods (2016–2018 and 2020–2022). Two annotators manually checked the corresponding abstracts.
@@
The first step in this study was to obtain metadata from the bioRxiv preprint platform using the official bioRxiv API [16]. The API has the following structure: server indicates the source of the data (e.g., bioRxiv or MedRxiv), interval indicates the desired date range, cursor indicates the pagination of the results, and format indicates the output format (e.g., JSON). The following data fields were collected through the API, and the retrieved metadata fields were mapped to the following structure in the dataset (see Table 2).
@@
This filtering aims to track changes between the first and last versions of the preprints, as well as changes that occurred when they were published in journals.
@@
For published preprints, the journal DOI is stored in the bioRxiv dataset. About two weeks after journal publication, bioRxiv automatically adds a link to the published article [17]. Detailed metadata for these articles was retrieved via the official Crossref API [18], by replacing the doi field with the Digital Object Identifier of the corresponding article. This query returns an extensive metadata package for the published version. The following fields were retrieved from the Crossref database (see Table 3).
@@
Title crossref_title Official title of the published version
@@
Published-online crossref_online_publication_date Date the article was first published online
@@
The official publication dates of 500 published articles were manually compared with the online_publication_date and issue_online_date obtained from Crossref. In most cases, the online_publication_date matched with the actual publication date of the article.
@@
In the third step, the dataset was categorized based on whether the preprints had DOIs linked to published articles on the bioRxiv preprint server.
@@
    If a published DOI was available for the preprint, the preprint was included in the published group.
@@
Although bioRxiv metadata is expected to include DOIs for all published preprints, verification revealed that in some cases the DOI of the published version is missing, even when the preprint has been published in a journal. For example, the fourth version of a preprint with DOI 10.1101/170381 was submitted to bioRxiv on August 6, 2020, but the same article was later published in Brain Structure and Function with DOI 10.1007/s00429-020-02136-0 on September 18, 2020. Nevertheless, the preprint still appears as preprint only in the bioRxiv record.
Previous studies have also documented this issue [7, 1, 5]. For example, Abdill and Blekhman (2019) found that 37.5% of 120 bioRxiv preprints were missing publication links. Fraser et al. (2021) reported that 7.6% of 12,788 preprints had missing publication links. Cabanac et al. (2021) reported that 60.3% of preprints marked as published on medRxiv and three other servers lacked corresponding publication links.
To identify such cases, we compared the titles of works marked as preprint only with the titles of published articles in the Crossref database. First, a separate request was sent to the Crossref database for each DOI, and matching titles were identified. Python’s SequenceMatcher library was then used to assess title similarity. The algorithm determines the longest common subsequence between two texts, accounts for the total length of matching parts, and calculates a similarity coefficient ranging from 0 to 1. A coefficient close to 1 indicates a complete match, whereas a value near 0 indicates substantial differences. Matches with a similarity index of $0.75$ or higher were classified into the Gray Zone category. This threshold was chosen because, in tests comparing preprints already marked as published in bioRxiv with their corresponding journal versions, results in this range proved most reliable in Figure 2.
Figure 2: Distribution of published papers by title match score
@@
In the final stage, several additional calculations were performed to enrich the dataset with further features. For example, the difference between the submission dates of the first and last versions of each preprint was calculated, as well as the difference between the official journal publication date and the last submission date on bioRxiv for published preprints. These values were stored in the custom_biorxivVersion_dateDifference and custom_submission&publication_dateDiff columns.
@@
The dataset is based on data collected from the bioRxiv [16] and Crossref [18] databases. It is stored in CSV format and is publicly available via Zenodo. The dataset consists of two main files:
@@
    • custom_status: Classification indicating the publication status of preprints: Published (officially published or confirmed with a DOI in bioRxiv), Preprint Only (unpublished), or Gray Zone (no official link in bioRxiv, but a possible publication identified based on title and author match).
    • biorxiv_published_doi: DOI of published versions of the preprint. This information is collected from two sources: (1) official publication DOIs provided by bioRxiv; (2) DOIs additionally identified in Gray Zone cases.
    • custom_submission&publication_dateDiff: The difference in days between the last preprint submission date and the publication date of the published version.
@@
    • author_count_diff: The difference between the number of authors in the preprint and the number of authors in the published version. Positive values indicate that the published version has more authors, while negative values indicate fewer.
    • title_match_score: The similarity score between the preprint and the published article titles. Values range from 0 to 1 and are calculated with the SequenceMatcher algorithm. Only matches within the 0.75–1.0 range were considered potential publication candidates.
    • author_match_score: Similarity score between the preprint and published article author lists. Ranges from 0 to 1 and is primarily used to verify Gray Zone cases, with higher values indicating stronger similarity.
@@
    • suspected_published_doi: Suspected DOI of the published article. Includes DOIs identified for Gray Zone preprints where a potential published version was found but could not be fully confirmed.
@@
Figure 3: (a) Combined annotations versus author match score for records characterized by the title match score equal to 0.75. Symbols are data points, and the curve comes from the fit to the logistic regression model (1). Data points are artificially jittered around TRUE and FALSE values to avoid overlap. (b) The fraction of preprints that were eventually published as journal articles versus the year of preprint submission. Triangles represent data without the gray zone records, circles take into account gray zone records, with the author match score larger than 0.47 used to distinguish between preprints and eventually published journal papers.
@@
where $\pi$ is interpreted as the probability of the dependent variable (annotation) equaling a success, $x$ reflects author match score and $\beta_{0},\beta_{1}$ are regression coefficients. The model was fitted on the set of non-NA data, resulting in $\beta_{0}=-2.99\pm 0.46$ and $\beta_{1}=6.36\pm 0.67$ (see the curve in Figure 3a). The location parameter $\mu=-\beta_{0}/\beta_{1}$ – the midpoint of the curve, where $pi$ crosses 1/2 is $\mu=0.47\pm 0.12$ where the standard deviation of $\mu$ has been calculated using propagation of uncertainty[22], i.e., $\sigma^{2}_{\mu}=|\beta_{0}/\beta_{1}|\sqrt{\sigma^{2}_{\beta_{0}}/\beta^{2}_{0}+\sigma^{2}_{\beta_{1}}/\beta^{2}_{1}}$. To assess the validity of the logistic regression approach, we performed 10-fold cross-validation[21], obtaining the accuracy=0.92 and Cohen’s $\kappa=0.81$.
Let us now see the ramifications of the author match score threshold $\mu$ on the possible use of the dataset. Figure 3b presents the fraction of preprints that were eventually published as journal articles $f$ versus the year of preprint submission. The first series, marked by triangles, disregards the gray zone papers, taking into account only records marked as “published”, which leads to a spectacular decrease in $f$ by over 30% between years 2016 and 2022. However, if we take into account that all gray zone records with author match score exceeding $\mu=0.47$ can be regarded as published ones, the decrease does not exceed 11%.
@@
PreprintToPaper dataset can be used in a variety of studies: one of the principal ideas is connected to the issue of information overload[23] in academia – a notion that is widely understood[24] (e.g., paper tsunami[25]) but for which no measures have been defined yet. The very construction of our dataset (i.e., a set of preprints submitted before / during the outbreak of the COVID-19 pandemic), along with the information about the period between the preprint submission, can help find the patterns of preprint-to-paper transformations under the stress exerted by the pandemic[25]. In the same manner, the question of the characteristics of “unpublished preprints” – treated in a similar way as uncited papers[26, 27] – could be examined using the provided dataset. These include stylometric factors: title and abstract length, text complexity (e.g., Herdan’s measure[28]), readability scores (e.g., Gunning FOG index[29]) as well as text sentiment, number of authors, and other that have been shown to play role in the previous studies linking textual factors and paper popularity[30, 31, 32].
@@
All custom scripts used in this study are publicly available in the GitHub repository: GitHub
@@
The code consists of scripts for different stages: downloading metadata from the bioRxiv API (download_biorxiv.py), selecting the first and last version and creating the pivot structure (filter_1st_and_last_version.py, pivot_biorxiv_data.py), getting published metadata from the Crossref API (download_crossref.py), categorizing preprints (categorize_preprints.py), detecting “Gray Zone” cases (find_missing_links_crossref.py), verifying by author match (verify_missing_links_by_author_match.py), standardizing dates and calculating differences (standardize_crossref_dates.py, merge_publication_dates.py, calculate_date_difference.py).

### v2

@@ Abstract
The PreprintToPaper dataset connects bioRxiv preprints with their corresponding journal publications, enabling large-scale analysis of the preprint-to-publication process. It comprises metadata for 145,517 preprints from two periods, 2016-2018 (pre-pandemic) and 2020-2022 (pandemic), retrieved via the bioRxiv and Crossref APIs. We selected the two periods to capture preprint–publication dynamics before and during the COVID-19 pandemic while avoiding transitional years. Each record includes bibliographic information such as titles, abstracts, authors, institutions, submission dates, licenses, and subject categories, alongside enriched publication metadata including journal names, publication dates, author lists, and further information. In addition to the main dataset, a version-history subset provides all available versions of preprints within the two selected periods, enabling analysis of how preprints evolve over time. Preprints are categorized into three groups: Published (formally linked to a journal article), Preprint Only (posted on a preprint server), and Gray Zone (potentially published in a journal but unlinked). To enhance reliability, title and author similarity scores were computed, and a human-annotated subset of 299 records was created to evaluate Gray Zone cases. The dataset supports diverse applications, including studies of scholarly communication, open science policies, bibliometric tool development, and natural language processing research on textual changes between preprints and the corresponding journal articles. The dataset is publicly available in CSV format via Zenodo*.
**footnotetext:
@@
The PreprintToPaper dataset [3], which we describe in this paper, links bioRxiv preprints with their subsequent journal publications, allowing for large-scale analysis of the preprint-to-publication process. It includes metadata on more than 145,000 preprints from two distinct periods (2016–2018, the pre-pandemic period, and 2020–2022, the COVID-19 pandemic period), with information on titles, authors, abstracts, institutions, licenses, submission and publication dates, and subject categories. Preprints are categorized into “Published”, “Preprint Only”, and “Gray Zone” cases, based on automated linking procedures combining Digital Object Identifier (DOI) information, title similarity, and author similarity. In addition to the main dataset, two subset files are provided: (i) a human-annotated subset of Gray Zone cases is provided to support evaluation of the automated matching approach (this subset includes records with title similarity (title_match_score = 0.75) and was independently checked by two annotators; further details on the annotation procedure are provided in the Methods section - Step 5: Gray Zone Verification) and (ii) a version-history subset that lists all available versions for preprints that have version one and at least one additional version. This subset is intended for detailed, version-level analyses of preprint evolution.
The motivation for creating the PreprintToPaper dataset is to provide a comprehensive resource for studying how preprints evolve into published journal articles or remain as preprints posted only on a preprint server [4, 5]. While some preprints are eventually linked to formal publications, many others stay as preprints without ever appearing in journals. This dataset is the systematic effort to automatically collect and link metadata from bioRxiv [6] with publication records, making it possible to analyze patterns in the preprint-to-publication process at scale [7]. We selected the 2016–2018 (pre-pandemic) and 2020–2022 (pandemic) periods to capture preprint–publication dynamics before and during the COVID-19 pandemic while avoiding transitional years.
@@
  • Natural language processing and text analysis: Supports research on linguistic and structural changes between preprints and the corresponding journal articles, such as title reformulations, abstract rewriting, and shifts in author order or contribution statements.
@@
Research on bioRxiv has shown that articles posted as preprints tend to receive more citations and online attention than comparable articles not deposited, although the causal mechanisms behind this effect remain uncertain [12]. Surveys of authors indicate that the main motivations for posting preprints are to increase visibility and accelerate dissemination, while concerns about peer review and lack of awareness are the main barriers [2]. Large-scale bibliometric analyses further confirm a citation and altmetric advantage for bioRxiv-linked publications, with preprints themselves increasingly cited and shared on social media [12]. During the COVID-19 pandemic, preprints on bioRxiv and medRxiv played an unprecedented role, being accessed, cited, and shared at higher rates, and influencing both journalistic reporting and policy discussions [1]. An analysis of biomedical publishing during the first months of the COVID-19 pandemic reveals a sharp rise in publication volume, faster acceptance times for COVID-19 papers, and reduced international collaboration, with these shifts occurring largely at the expense of non-COVID-19 research [8]. Other studies of bioRxiv preprints have found that approximately 30% remain as preprints posted only on a preprint server, that many are posted too late to enable substantive feedback, and that nearly half of preprints published in a journal are concentrated among four major publishers, suggesting that the platform is used more for priority setting and visibility than for pre-publication peer review [4, 13].
@@
Europe PMC[16] maintains a large corpus of preprints and journal articles, including bioRxiv preprints, and links them to their corresponding publications.
Compared with the three related preprint resources discussed above, the PreprintToPaper dataset offers two main advantages:
  • It stores both the initial and final versions of bioRxiv preprints, along with detailed metadata (e.g., title, authors, journal, and publication date) for the corresponding journal articles. In addition, a separate version-history file contains all available preprint versions for cases with more than one version within our study periods.
  • It identifies articles that are not formally linked but may have been published in a journal (Gray Zone cases) by using title and author similarity, which tolerate minor title changes and variations in author order, thereby going beyond exact-match linking.
Find a detailed comparison of the four datasets in Table 1.
Table 1: Comparison of the covid19_preprints, Rxivist, Europe PMC, and PreprintToPaper datasets
Aspect covid19_preprints Rxivist Europe PMC corpus PreprintToPaper (Ours)
Time span 2017-2021 Nov 2013 - Feb 2023 2013 - 2023 2016-2018 (pre-pandemic) and 2020-2022 (pandemic)
Coverage Multiple preprint servers (bioRxiv, arXiv, OSF, ESSOAR, Figshare, Zenodo, etc.) bioRxiv and medRxiv Multiple 32 life-science preprint servers bioRxiv
Scope COVID-19 related preprints only All subjects (bioRxiv + medRxiv) All subjects (32 servers) All subjects (bioRxiv)
Size 66,130 records 223,541 records 761,123 records 145,517 records
Preprint information Source, identifier (DOI), posted date, title, abstract Title, abstract, DOI, category, submission date, authors, institutions Title, abstract, DOI, submission date, authors, full text of COVID-19 and Europe PMC funder preprints in JATS XML, author affiliations, license, funding, versions, status ( Withdrawal/removal ) Title, abstract (initial + last), DOI, category, submission date (initial + last), authors (initial + last), corresponding authors, institutions, license, version number, category, fulltext JATS XML
Publication information None Published journal article DOI, journal name Published journal article DOI, journal name, publication date, citations, references Published journal article DOI, title, authors, journal name, publication date (from Crossref)
Versioning Only one version per preprint Only one version per preprint All versions (except from bioRxiv, medRxiv, SSRN) All versions at the relevant period (bioRxiv)
Publication linking None Published journal article DOIs partially incomplete (crawling stopped early) Official bioRxiv links enriched with journal metadata + an investigation of unlinked preprints where an exact match of title and first author is required. Even minor changes in title or author names are ignored; as a result, many preprints with slightly different titles are not linked to their journal papers Official bioRxiv links enriched with Crossref + additional Gray Zone investigation where both titles and full author lists are compared using similarity scores. Preprint–paper pairs with title similarity between 75% and 100% are identified and added as potential links ($\approx$19,000)
Two further resources provide data related to preprint-publication links: PreprintMatch [17] and the Crossref relationships dataset [18]. In contrast to PreprintToPaper, they focus on DOI-to-DOI linkage and matching infrastructure, rather than delivering a bioRxiv-specific dataset with enriched metadata.
@@ Methods
Figure 1 illustrates the data generation workflow for the PreprintToPaper dataset. The process consists of six main steps: (1) collecting metadata from the bioRxiv preprint platform; (2) retrieving additional metadata from Crossref based on the published journal article DOIs listed in the bioRxiv records; (3) categorizing preprints into published and preprint-only; (4) identifying unlinked publications in bioRxiv (Gray Zone); (5) verifying Gray Zone preprints using author-based metrics (author match score and author count difference); and (6) generating the final dataset.
Figure 1: PreprintToPaper dataset creation workflow. Final enriched dataset and two derived subsets are generated: (i) a human-annotated Gray Zone subset of 299 title-matched cases (title match score = 0.75) and (ii) a version history subset, in which multi-version preprints that retain version 1 and at least one later version within the study periods are included.
@@
The first step in this study was to obtain metadata from the bioRxiv preprint platform using the official bioRxiv API (https://api.biorxiv.org/details/[server]/[interval]/[cursor]/[format]. The API has the following structure: server indicates the source of the data (e.g., bioRxiv or MedRxiv), interval indicates the desired date range, cursor indicates the pagination of the results, and format indicates the output format (e.g., JSON). The following data fields were collected through the API, and the retrieved metadata fields were mapped to the following structure in the dataset (see Table 2).
@@
In addition to the main table, a separate version-history subset was created that contains all available versions for preprints with a first version and at least one subsequent version within our study periods. The filtering applied to the main dataset is designed to track changes between the first and last preprint versions, as well as changes that occur upon journal publication, whereas the version-history subset supports more detailed, version-level analyses. However, because the bioRxiv API was observed to return identical abstract text across different versions of the same preprint, abstracts were not included in the version-history subset.
@@
For preprints that were later published as journal articles, the corresponding journal DOI is stored in the bioRxiv dataset. About two weeks after journal publication, bioRxiv automatically adds a link to the corresponding journal article (https://www.biorxiv.org/about/FAQ). Detailed metadata for these articles was retrieved via the official Crossref API (https://api.crossref.org/works/[doi]), by replacing the doi field with the DOI of the corresponding article. This query returns an extensive metadata package for the corresponding journal article. The following fields were retrieved from the Crossref database (see Table 3).
@@
Title crossref_title Official title of the corresponding journal article
@@
Published-online crossref_online_publication_date Date the journal article was first made available online
@@
The official publication dates of 500 journal articles were manually compared with the online_publication_date and issue_online_date obtained from Crossref. In most cases, the online_publication_date matched with the actual publication date of the article.
@@
In the third step, the dataset was categorized based on whether the preprints had DOIs linked to journal articles on the bioRxiv preprint server.
@@
    If a journal article DOI was available for the preprint, the preprint was included in the published group.
@@
Although bioRxiv metadata is expected to include DOIs for all preprints that have been published as journal articles, verification revealed that in some cases the DOI of the corresponding journal article is missing, even when the preprint has been published in a journal. For example, the fourth version of a preprint with DOI 10.1101/170381 was submitted to bioRxiv on August 6, 2020, but the same article was later published in Brain Structure and Function with DOI 10.1007/s00429-020-02136-0 on September 18, 2020. Nevertheless, the preprint still appears as preprint only in the bioRxiv record.
Previous studies have also documented this issue [7, 1, 5]. For example, Abdill and Blekhman (2019) found that 37.5% of 120 bioRxiv preprints were missing publication links. Fraser et al. (2021) reported that 7.6% of 12,788 preprints had missing publication links. Cabanac et al. (2021) reported that 60.3% of preprints marked as journal-published on medRxiv and three other servers lacked corresponding publication links.
To identify such cases, we compared the titles of works marked as preprint only with the titles of the corresponding journal articles in the Crossref database. First, a separate request was sent to the Crossref database for each DOI, and matching titles were identified. Python’s SequenceMatcher library was then used to assess title similarity. The algorithm determines the longest common subsequence between two texts, accounts for the total length of matching parts, and calculates a similarity coefficient ranging from 0 to 1. A coefficient close to 1 indicates a complete match, whereas a value near 0 indicates substantial differences.
To assess our choice of similarity measure, we compared Python’s SequenceMatcher with two FuzzyWuzzy metrics on a random sample of 100 preprint and publication title pairs from our dataset, all of which were confirmed as published in a journal. In 58% of the cases, all three measures produced identical scores, and in the remaining cases the differences were small and did not change the classification of any pair with respect to our 0.75 threshold. We therefore retained SequenceMatcher as our main measure, as it is part of the Python standard library.
Matches with a similarity index of $0.75$ or higher were classified into the Gray Zone category. This threshold was chosen because, in tests comparing preprints already marked as published in bioRxiv with their corresponding journal versions, results in this range proved most reliable in Figure 2.
Figure 2: Distribution of journal articles by title match score.
@@
In the final stage, several additional calculations were performed to enrich the dataset with further features. For example, the difference between the submission dates of the first and last versions of each preprint was calculated, as well as the difference between the official journal publication date and the last submission date on bioRxiv for preprints published in journal. These values were stored in the custom_biorxivVersion_dateDifference and custom_submission&publication_dateDiff columns.
@@
The dataset is based on data collected from the bioRxiv and Crossref databases. It is stored in CSV format and is publicly available via Zenodo [3] (https://doi.org/10.5281/zenodo.17992421). The dataset consists of three main files:
@@
    • custom_status: Classification indicating the publication status of preprints: Published (officially published in a journal and confirmed with a DOI in bioRxiv), Preprint Only (posted on a preprint server), or Gray Zone (no official link in bioRxiv, but a possible publication identified based on title and author match).
    • biorxiv_published_doi: DOI of the corresponding journal article. This information is collected from two sources: (1) official publication DOIs provided by bioRxiv; (2) DOIs additionally identified in Gray Zone cases.
    • custom_submission&publication_dateDiff: The difference in days between the last preprint submission date and the publication date of the corresponding journal article.
@@
    • author_count_diff: The difference between the number of authors in the preprint and the number of authors in the corresponding journal article. Positive values indicate that the preprints published in journal have more authors, while negative values indicate fewer.
    • title_match_score: The similarity score between the preprint and the journal article titles. Values range from 0 to 1 and are calculated with the SequenceMatcher algorithm. Only matches within the 0.75–1.0 range were considered potential publication candidates.
    • author_match_score: Similarity score between the preprint and corresponding journal article author lists. Ranges from 0 to 1 and is primarily used to verify Gray Zone cases, with higher values indicating stronger similarity.
@@
    The Version History Subset is a separate, version-level file that contains all available versions for preprints that have version one and at least one additional version within our study periods (2016–2018 and 2020–2022). Preprints with only a single version or without the first version present are not included in this file. Each row corresponds to a bioRxiv version and includes the DOI, version number, submission date, title, authors, category, license, and other core bioRxiv metadata. This subset allows detailed analyses of how preprints evolve across multiple versions.
  3. 3.
@@
    • suspected_published_doi: Suspected DOI of the journal-published article. Includes DOIs identified for Gray Zone preprints where a potential corresponding journal article was found but could not be fully confirmed.
@@
Figure 3: (a) Combined annotations versus author match score for records characterized by the title match score equal to 0.75 (see text for rationale). Symbols are data points, and the curve comes from the fit to the logistic regression model (1). Data points are artificially jittered around TRUE and FALSE values to avoid overlap. (b) The fraction of preprints that were eventually published as journal articles versus the year of preprint submission. Triangles represent data without the Gray Zone records; circles take into account Gray Zone records, with an author match score larger than 0.47 (suggested by logistic regression analysis; see text in Technical Validation) used to distinguish between preprints and journal-published articles.
@@
where $\pi$ is interpreted as the probability of the dependent variable (annotation) equaling a success, $x$ reflects author match score and $\beta_{0},\beta_{1}$ are regression coefficients. The model was fitted on the set of non-NA data, resulting in $\beta_{0}=-2.99\pm 0.46$ and $\beta_{1}=6.36\pm 0.67$ (see the curve in Figure 3a). The location parameter $\mu=-\beta_{0}/\beta_{1}$ – the midpoint of the curve, where $\pi$ crosses 1/2 is $\mu=0.47\pm 0.12$ where the standard deviation of $\mu$ has been calculated using propagation of uncertainty[22], i.e., $\sigma^{2}_{\mu}=|\beta_{0}/\beta_{1}|\sqrt{\sigma^{2}_{\beta_{0}}/\beta^{2}_{0}+\sigma^{2}_{\beta_{1}}/\beta^{2}_{1}}$. To assess the validity of the logistic regression approach, we performed 10-fold cross-validation[21], obtaining the accuracy=0.92 and Cohen’s $\kappa=0.81$.
Let us now see the ramifications of the author match score threshold $\mu$ on the possible use of the dataset. Figure 3b presents the fraction of preprints that were eventually published as journal articles $f$ versus the year of preprint submission. The first series, marked by triangles, disregards the gray zone papers, taking into account only records marked as “published”, which leads to a spectacular decrease in $f$ by over 30% between years 2016 and 2022. However, if we take into account that all gray zone records with author match score exceeding $\mu=0.47$ (obtained as threshold from the logistic regression analysis) can be regarded as journal-published ones, the decrease does not exceed 11%.
@@
PreprintToPaper dataset can be used in a variety of studies: one of the principal ideas is connected to the issue of information overload[23] in academia – a notion that is widely understood[24] (e.g., paper tsunami[25]) but for which no measures have been defined yet. The very construction of our dataset (i.e., a set of preprints submitted before / during the outbreak of the COVID-19 pandemic), along with the information about the period between the preprint submission, can help find the patterns of preprint-to-paper transformations under the stress exerted by the pandemic[25]. Similarly, the characteristics of “preprints posted on preprint server” – treated in a manner analogous to uncited papers[26, 27] – can be examined using the provided dataset. These include stylometric factors: title and abstract length, text complexity (e.g., Herdan’s measure[28]), readability scores (e.g., Gunning FOG index[29]) as well as text sentiment, number of authors, and other that have been shown to play role in the previous studies linking textual factors and paper popularity[30, 31, 32].
Data Availability
The dataset supporting this study has been stored in Zenodo and is openly accessible at https://doi.org/10.5281/zenodo.17992421. Further details about the dataset contents are provided in the Data Record section of this manuscript.
@@
All custom scripts used in this study are publicly available in the GitHub repository: GitHub (https://github.com/badalovafidan/preprint_to_paper_dataset)
@@
The code consists of scripts for different stages: downloading metadata from the bioRxiv API (download_biorxiv.py), selecting the first and last version and creating the pivot structure (filter_1st_and_last_version.py, pivot_biorxiv_data.py), creating additional version-history subset (create_version_history.py), getting the corresponding journal article metadata from the Crossref API (download_crossref.py), categorizing preprints (categorize_preprints.py), detecting “Gray Zone” cases (find_missing_links_crossref.py), verifying by author match (verify_missing_links_by_author_match.py), standardizing dates and calculating differences (standardize_crossref_dates.py, merge_publication_dates.py, calculate_date_difference.py).